OpenStudy (anonymous):

cos^6A-sin^6A=(cos^2A-sin^2A)*(1-sin^2A*cos^2A)

OpenStudy (anonymous):

need help in breaking in simpler terms

OpenStudy (anonymous):

power 6 ?? -_-

OpenStudy (ajprincess):

\(\cos^6A-\sin^6A=(\cos^2A)^3-(\sin^2A)^3\) \((a^3-b^3)=(a-b)(a^2+ab+b^2)\) here \(a=\cos^2A\) \(b=\sin^2A\) also \(\cos^2A+\sin^2A=1\) Does that help? @koli123able

OpenStudy (anonymous):

thanks once again!! :-)

OpenStudy (ajprincess):

Welcome:)

OpenStudy (anonymous):

(cos^2A-sin^2A)(cos^4A+sin^4A+cos^2A*sin^2A)

OpenStudy (anonymous):

what to do after this ^^

OpenStudy (anonymous):

@ajprincess

OpenStudy (zehanz):

How many of these probs you've got? Looks like you can fill the entire sunday with it :/

OpenStudy (anonymous):

well I did almost just this one is left I did 53 questions since morning

OpenStudy (zehanz):

Wow! Hats off to you!

OpenStudy (anonymous):

can u help me with this last one

OpenStudy (anonymous):

help plzz

OpenStudy (anonymous):

????

OpenStudy (zehanz):

OK, I'll give it a try. Give me a few secs :)

OpenStudy (ajprincess):

\((cos^2A-sin^2A)(cos^4A+sin^4A+cos^2A*sin^2A)\) \(=(cos^2A-sin^2A)((cos^2A)^2+(sin^2A)^2+cos^2A+sin^2A)\) \((cos^2A)^2+(sin^2A)^2=1\)

OpenStudy (zehanz):

Shouldn't that "+" at the end of the middle line be a dot?

OpenStudy (ajprincess):

yup u r right:)@ZeHanz. really sorry abt that typo

OpenStudy (zehanz):

We have so far: \((\cos^2A-\sin^2A)(\cos^4A+\sin^4A+\cos^2A \cdot \sin^2A)\). By peeking at the RHS, we see that the first factor is there too. We just need to do something with the second one: \(\cos^4A+\sin^4A+\cos^2A \cdot \sin^2A\)= \(\cos^4A+\sin^4A+\cos^2A(1-\cos^2A)\)= \(\cos^4A+\sin^4A+\cos^2A-\cos^4A\)= \(\sin^4A+\cos^2A\)= \((1-\cos^2A)^2+\cos^2A\)= \(1-2\cos^2A+\cos^4A+\cos^2A\)= \(1-\cos^2A+\cos^4A\)= \(1-\cos^2A(1-\cos^2A)\)= \(1-\cos^2A\sin^2A\) So if we multiply this with \(\cos^2A-\sin^2A\), we're done!